The radii of curvature of a double convex lens are 15cm and 30cm and its refractive index is 1.5. then its focal length is
      A. +10  
  1.  -10 
  1. +20
  1. -20

Grade:12

2 Answers

CHALLA NAVNEESWAR REDDY
7 years ago
we know the formula
1/f=R.I(-1)(1/R1-1/R2)
=1.5(-1)(1/15-1/30)  ,     =-1.5(1/30)
=      1/f=  -1/20 ,     f=-20
focal length=-20cm
Dhruv Shah
7 years ago
Answer is +20 bcoz one has radius of curvature +15 and other has -30 and so on subsituting we get (Nrel-1)(1/r1-1/r2)(3/2-1)(1/15-(-1/30))(1/2)(1/10)1/20F=20cm